OpenStudy (anonymous):

on each of the days monday through friday , toni spent hour commuting to work and hour commuting back home . what fraction of the total number of hours in these five days did she spend commuting !! 1/12 1/24 5/12 5/24 10/168

OpenStudy (kropot72):

The times for one day are enough to solve this. Toni spends 2 hours commuting every 24 hours. Fraction of 24 hours spent commuting =\[\frac{2}{24}\] One of the choices equals this fraction. Which one is it?

OpenStudy (kropot72):

\[\frac{2}{24}=\frac{1}{12}\]when we divide the numerator and the denominator of the first fraction by 2.

OpenStudy (anonymous):

5/12

OpenStudy (kropot72):

The correct choice is 1/12

OpenStudy (unklerhaukus):

ie one twelfth of tonis time during the week in spent commuting

OpenStudy (anonymous):

but there r 5 days means 5time cus one day she spend 2hours means 10hours per week !!? right

OpenStudy (anonymous):

ohhh ok the fraction I got it ok thank you gys sorry .....^ ^alooha

OpenStudy (unklerhaukus):

\[\frac {5\times2}{5\times24}=\frac{10}{120}=\frac{1}{12}\]